40x10^5 grams of Na2CO3= How many particles of Na2CO3
1
Expert's answer
2019-05-17T02:43:12-0400

"n(Na2CO3)=\dfrac{40*10^5g}{106g/mol}=37735.85moles"

"N(Na2CO3)=37735.85moles*6.02*10^{23}=2.272*10^{28}(particles)"